OpenStudy (anonymous):

How will adding the value 1000 affect the mean and median of the data set 5, 10, 17, 19, 20? A. The mean and the median increase by the same amount. B. The median increases and the mean stays the same. C. The mean increases and the median stays the same. D. The mean increases more than the median increases.

OpenStudy (anonymous):

okay, do you know what the mean and median is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

so 17 plus 19/2=

OpenStudy (anonymous):

26.5

OpenStudy (anonymous):

17 plus 19 is 36

OpenStudy (anonymous):

ya i now that

OpenStudy (anonymous):

divide that by 2 and get 18

OpenStudy (anonymous):

how did you come up with 26.5

OpenStudy (anonymous):

i did 17+19/2=

OpenStudy (anonymous):

okay let me rewrite that 17+19=? divide that answer by 2

OpenStudy (anonymous):

ok and it = 18

OpenStudy (anonymous):

okay without the 1000 the median is 17...so he median increased by one

OpenStudy (anonymous):

which means we can eliminate C

OpenStudy (anonymous):

and we know the mean will not stay the same because we are adding another huge number

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which leaves us with A and D

OpenStudy (anonymous):

And just looking at the set of numbers we know the mean will not increase by one (because the median increased by one)

OpenStudy (anonymous):

Which gives you?

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yaaaaaaay!!!!! thx

OpenStudy (anonymous):

You did a great job as well

OpenStudy (anonymous):

medal for you and thx

OpenStudy (anonymous):

and one for you for not rushing the answer :)

OpenStudy (anonymous):

thx